OpenStudy (anonymous):

Simplify. Show your work. 5 1/3 +(-3 9/18)

jhonyy9 (jhonyy9):

first 9/18 = 1/2 ok ?

jhonyy9 (jhonyy9):

than 5 1/3 = (5*3 +1)/3 = 16/3 can you continue it ?

OpenStudy (anonymous):

nop i never learn how to that

OpenStudy (anonymous):

well 16/3 is 5.33 how i put i a fracion

jhonyy9 (jhonyy9):

and now you need to transform the second fraction in form of the first can you make it ?

jhonyy9 (jhonyy9):

-3 1/2 =

jhonyy9 (jhonyy9):

inside parantheses there are -3 9/18 yes ?

jhonyy9 (jhonyy9):

so but you can simplifie there by 9 and will get -3 1/2 ok ?

OpenStudy (anonymous):

5 1/3 = 16 /3

OpenStudy (anonymous):

-3 9/18= -45/18

jhonyy9 (jhonyy9):

yes than -3 1/2 = how many will be ?

jhonyy9 (jhonyy9):

courage please

OpenStudy (anonymous):

16/3 + -45/18 and what i need to do next

jhonyy9 (jhonyy9):

wha -45/18 you can working more easy with 1/2 nope ?

OpenStudy (anonymous):

so i need to change -45 /18 to 1/2

jhonyy9 (jhonyy9):

no there are -3 1/2 =

OpenStudy (anonymous):

what happend if i work w -45/18

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

hey

jhonyy9 (jhonyy9):

not is right

OpenStudy (anonymous):

16/3 + -45/18 = -29/6

OpenStudy (anonymous):

iam wrong

OpenStudy (anonymous):

or -29/18

OpenStudy (anonymous):

i am right or not

OpenStudy (anonymous):

5 1/3 +(-3 9/18) = 5 1/3 - 3 9/18 = 16/3 - 63/18 = 32/6 - 21/6 = 11/6 = 1 5/6

jhonyy9 (jhonyy9):

5 1/3 = 16/3 - 3 9/18 = -3 1/2 = - 7/2 16/3 - 7/2 = (2*16 - 7*3)/6 =(32-21)/6 = 11/6 yes you are right

jhonyy9 (jhonyy9):

good luck bye

OpenStudy (anonymous):

I am still lost how to do this here. Please help me understand